# Physics-aware Machine Learning Revolutionizes Scientific Paradigm for Machine Learning and Process-based Hydrology

### v1

@@ 3.2.1 Physics in Hydrodynamic modeling
1D models. 1D hydrodynamic models have been widely used in modeling water flows [284]. Many hydraulic situations can make the 1D assumption, either because a more detailed solution is unnecessary (e.g., the purpose does not require knowledge in other dimensions) or because the flow is markedly 1D, such as in a confined channel or in a channel flow. Generally, 1D models solve equations derived by ensuring mass and momentum conservation, which yields the well-known 1D Saint-Venant equations,
@@
where $x$ is the distance along the river channel; $t$ is time; $q$ is the water inflow per unit length of the channel from land surface and subsurface runoff, groundwater, and precipitation; $u$ and $h$ are the dynamics of velocity and water depth along the river channel, respectively; $g$ is gravity; $S$ is riverbed slope; and $S_{f}$ is the friction slope, which can be calculated using the Chezy–Manning equation [285].
2D models. The 2D models represent water or flood flow as a two-dimensional field with the assumption that the third dimension, water depth, is shallow in comparison to the other two dimensions [286, 287]. Most approaches solve two-dimensional shallow water equations, which represent mass and momentum conservation in a plane, and can be obtained by depth-averaging the Navier-Stokes equations,
@@
where $x$ and $y$ are the two Cartesian directions; $q_{x}$ and $q_{y}$ are the $x$ and $y$ components of the discharge per unit width vector $\mathbf{q}$; $u$ and $v$ are the $x$ and $y$ components of the flow velocity, respectively; $z$ is the bed elevation; and $n$ is the Manning’s friction coefficient. In flood events, the key simplification introduced by the local inertial approximation is based on the assumption that the convective acceleration terms are negligible compared with the other terms and, therefore, can be ignored.
3D models. For many scales of water flow, complex 3D representation of flow dynamics has tended to be regarded as unnecessary, as a 2D shallow water approximation may be adequate [288]. However, modeling vertical turbulence, vortices, and floods occurring due to dam breaks is important. 3D models are therefore developed to allow the representation of vertical features. Some of these models solve the horizontal flow with 2D shallow water equations and include a quasi-3D extension to model velocity in vertical layers. Other 3D models are derived from the 3D Navier-Stokes equations, which describe the motion of fluid substances and are usually written as
$\begin{gathered}\frac{\partial u}{\partial t}+u\cdot\nabla u+\frac{1}{\rho}\nabla p=g+\mu\nabla\cdot\nabla u,\\ \text{ Incompressibility condition:}\quad\nabla\cdot u=0,\end{gathered}$ (5)
where $u$ is the velocity; $\rho$ is the fluid density; $p$ is pressure; $g$ is gravitational acceleration; and $\mu$ is kinematic viscosity. The incompressibility condition is based on the assumption that material density is constant within a fluid parcel. Depending on the representation of the processes, these models may be divided into two broad categories: Eulerian (grid-based) and Lagrangian (particle-based).
The following sections: Hydrodynamic Modeling Meets Physics-aware Machine Learning, Proposed HydroPML-based Flood Forecasting (Fig. 6), and Summary and Future Directions will be made available after the paper is accepted.
@@ Acknowledgements
The work is jointly supported by the German Federal Ministry of Education and Research (BMBF) in the framework of the international future AI lab "AI4EO – Artificial Intelligence for Earth Observation: Reasoning, Uncertainties, Ethics and Beyond" (grant number: 01DD20001) and the project Inno__\__MAUS (grant number: 02WEE1632B), by German Federal Ministry for Economic Affairs and Climate Action in the framework of the "national center of excellence ML4Earth" (grant number: 50EE2201C), by the Excellence Strategy of the Federal Government and the Länder through the TUM Innovation Network EarthCare and by Munich Center for Machine Learning.

### v4

@@ Appendix B. Physical Methods in Hydrodynamic Modeling
1D models. 1D hydrodynamic models have been widely used in modeling water flows [334]. Many hydraulic situations can make the 1D assumption, either because a more detailed solution is unnecessary (e.g., the application does not necessitate knowledge in other dimensions) or because the flow is distinctly 1D, such as in a confined channel or in a channelized flow. Typically, 1D models solve equations derived to ensure conservation of mass and momentum, resulting in the well-known 1D Saint-Venant equations,
@@
where $x$ denotes the distance along the river channel; $t$ denotes time; $q$ represents the water inflow per unit length of the channel from land surface and subsurface runoff, groundwater, and precipitation; $u$ and $h$ represent the dynamics of velocity and water depth along the river channel, respectively; $g$ represents gravity; $S$ represents the riverbed slope; and $S_{f}$ represents the friction slope, which can be computed using the Chezy–Manning equation [335].
2D models. The 2D models depict water or flood flow as a two-dimensional field with the assumption that the third dimension, water depth, is shallow compared to the other two dimensions [336, 337]. Most methods solve two-dimensional shallow water equations, which describe conservation of mass and momentum in a plane and are derived by depth-averaging the Navier-Stokes equations,
@@
here, $x$ and $y$ represent the two Cartesian directions; $q_{x}$ and $q_{y}$ denote the $x$ and $y$ components of the discharge per unit width vector $\mathbf{q}$; $u$ and $v$ are the $x$ and $y$ components of the flow velocity, respectively; $z$ denotes the bed elevation; and $n$ represents the Manning’s friction coefficient. In flood situations, the significant simplification introduced by the local inertial approximation relies on the assumption that the convective acceleration terms are negligible compared to other terms and can thus be disregarded.
3D models. For various water flow scenarios, there has been a tendency to consider the detailed representation of flow dynamics in 3D unnecessary, favoring the adequacy of a 2D shallow water approximation [338]. Nevertheless, it is crucial to model vertical turbulence, vortices, and floods caused by dam breaks. Consequently, 3D models have been developed to incorporate vertical features. Some of these models solve horizontal flow using 2D shallow water equations and incorporate a quasi-three-dimensional extension to represent velocity in vertical layers. Other 3D models are based on the 3D Navier-Stokes equations, which describe the movement of fluid substances and are typically formulated as,
$\begin{gathered}\frac{\partial u}{\partial t}+u\cdot\nabla u+\frac{1}{\rho}\nabla p=g+\mu\nabla\cdot\nabla u,\\ \quad\nabla\cdot u=0,\end{gathered}$ (5)
where $u$ represents velocity; $\rho$ denotes fluid density; $p$ stands for pressure; $g$ is gravitational acceleration; and $\mu$ represents kinematic viscosity. The incompressibility condition ($\quad\nabla\cdot u=0$) assumes that material density remains constant within a fluid parcel. Depending on how the processes are represented, these models can be categorized into two main types: Eulerian (grid-based) and Lagrangian (particle-based).
@@ Acknowledgements
The work is jointly supported by the Helmholtz Association through the Framework of the Helmholtz Excellent Professorship “Data Science in Earth Observation - Big Data Fusion for Urban Research”(grant number: W2-W3-100), by the German Federal Ministry of Education and Research (BMBF) in the framework of the international future AI lab "AI4EO – Artificial Intelligence for Earth Observation: Reasoning, Uncertainties, Ethics and Beyond" (grant number: 01DD20001) and by German Federal Ministry for Economic Affairs and Climate Action in the framework of the "national center of excellence ML4Earth" (grant number: 50EE2201C).